OpenStudy (anonymous):

Integration Problem Integral sqrt(1-x^2) dx from -1 to 1 Please give me the full solutions Thank in advanced.

sam (.sam.):

\[\int\limits \sqrt{1-x^2} \, dx?\]

OpenStudy (anonymous):

Yes from -1 to 1

sam (.sam.):

let x=sinu dx=cosu du

sam (.sam.):

\[\sqrt{1-x^2} =\sqrt{1-\sin ^2(u)}\text{ = }\cos (u)\text{ and }u=\sin ^{-1}(x)\]

sam (.sam.):

\[\int\limits \cos ^2(u) \, du\] Half angle formula \[\int\limits \left(\frac{1}{2} \cos (2 u)+\frac{1}{2}\right) \, du\] \[\text{}=\int\limits \frac{1}{2} \, du+\frac{1}{2}\int\limits \cos (2 u) \, du\] \[\text{For} ~\cos (2 u)\text{, substitute }k=2 u\text{ and }dk=2d u:\] \[\frac{1}{4}\int\limits \cos (k) \, dk+\int\limits \frac{1}{2} \, du\] \[\frac{\sin (k)}{4}+\frac{u}{2}\] \[\Huge [\frac{1}{2} \sqrt{1-x^2} x+\frac{1}{2} \sin ^{-1}(x)]_{-1}^{1}\]

OpenStudy (anonymous):

Can I ask you some thing , integral sqrt(1-x^2) dx from -1 to 1 = pi/2

sam (.sam.):

yes

sam (.sam.):

integral sqrt(1-x^2) dx from -1 to 1 = pi/2

OpenStudy (anonymous):

Oo thx

OpenStudy (anonymous):

U = sin^-1 (x) ???